OpenStudy (anonymous):

I have a continuous function f for all x in R satisfied \[\int_{-\infty}^{x} {f(u)du}\]\ for all x in R. \[\f>0\]\ for all x in R?

OpenStudy (anonymous):

What is your question exactly??

OpenStudy (turingtest):

let me get this straight...\[f:\forall x\in\mathbb R\]we have\[\int_{-\infty}^{x}f(u)du\]and you want to know if that means \[f>0:\forall x\in\mathbb R\]?

OpenStudy (anonymous):

my question : \[\forall x \in \mathbb{R}, f>0\]. True or false?

OpenStudy (turingtest):

are you told any more info, like that the integral converges or something?

OpenStudy (anonymous):

Sorry, the above integral >0 \[\forall x \in \mathbb{R}\]

OpenStudy (turingtest):

so we have the full question as: for\[f:\forall x\in\mathbb R\]we have\[\int_{-\infty}^{x}f(u)du>0\]does this mean that \[\forall x\in\mathbb R,~f>0\]?

OpenStudy (anonymous):

Additionally,f is continuous. Does \[f>0, \forall x \in \mathbb{R}\]?

OpenStudy (anonymous):

bookmark

OpenStudy (zarkon):

\(f\) does not have to be positive for all values of \(x\)

OpenStudy (anonymous):

Can you give me an example?

OpenStudy (anonymous):

Can we take the derivative of this inequality with respect to x: \[ \frac{d}{dx} (\int_{-\infty}^{x} f(u)du) > \frac{d}{dx}(0) \]? Or is it unnecessary?

OpenStudy (anonymous):

\[\int_{\infty}^x(\sin(x)+\frac{1}{2})dx\] could do it

OpenStudy (zarkon):

that integral does not converge for ant value of x

OpenStudy (zarkon):

*any

OpenStudy (anonymous):

oh no it won't do it since it wont

OpenStudy (anonymous):

what zarkon said

OpenStudy (zarkon):

try making a function f that looks like this...